what is the y-coordinate of the center of mass of a plane of the same shape as the histogram of the distribution ?


1
Expert's answer
2021-06-03T17:02:13-0400

"C_{y}=\frac{\sum C_{iy}A_i}{\sum A_i}"

where "A_i=df_i" is area i-th part of histogram,

"C_{iy}=f_i\/2" is y-coordinate of the center of mass of i-th part of histogram,

"f_i" is frequency,

"d" is width of interval.


Then:

"C_y=\frac{\sum df_i^2}{2\sum df_i}"